# Theory-Grounded Evaluation of Human-Like Fallacy Patterns in LLM Reasoning

### v2

@@ 1 Introduction
Language models increasingly solve complex tasks [1, 2]. We ask a simple question: when they err on controlled reasoning problems, do their errors align with human fallacies? Human reasoning shows systematic, repeatable fallacies [3, 4, 5, 6, 7]. These are not random mistakes; they follow predictable patterns across tasks. Understanding whether LLM errors share these patterns is useful both scientifically and for deployment in settings that require reliable reasoning [8, 9].
To rigorously investigate this question, we leverage the Erotetic Theory of Reasoning (ETR) [10, 11, 12], a formal cognitive theory that precisely predicts human reasoning patterns across multiple domains. ETR posits that human reasoning operates by maintaining disjunctive alternatives and filtering these alternatives when taking on new information, a process that can systematically lead to error. ETR provides formal specifications of both how and when humans will make specific reasoning errors, allowing us to generate arbitrary reasoning problems with predictable failure patterns.
ETR has an open-source implementation, PyETR [13]. Using PyETR, we mechanically generate reasoning problems predicted to elicit specific fallacies. We evaluate 38 models on a fixed corpus of 383 such problems.
Our experimental investigation evaluates 38 language models ranging from smaller models (e.g., Mistral 7B Instruct v0.1) to larger systems (e.g., GPT-4.5, Claude 3.7), treating Chatbot Arena Elo as one capability proxy [14]. We observe that as this proxy increases, a larger share of logically incorrect answers align with ETR-predicted human fallacies ($\rho=0.360,p=0.0265$). We do not claim causality and restrict interpretation to this evaluation setting.
While language models often improve on many benchmarks [1, 15], our results indicate that within our task domain, the proportion of errors matching ETR-predicted fallacies increases with Chatbot Arena Elo as a capability proxy. These findings are consistent with overlap between LLM error patterns and human fallacy patterns and do not require commitments about underlying cognitive mechanisms.
We check the robustness of our result by using metrics other than Chatbot Arena Elo where available for a subset of the 38 models. We found significant Spearman correlation with estimates of training compute [16] ($\rho=0.489,p=0.0334$) and significant exponential fit for mean score on the HELM capabilities benchmark [17] ($r=0.796,p=0.0103$).
@@
The remainder of the paper is organised as follows. Section 2 introduces ETR and PyETR. Section 3 details methodology. Section 4 presents results. Section 5 discusses implications.
@@
Substantial prior literature in psychology and cognitive science documents irrationalities in human judgement and decision-making [3, 4, 5, 18]. In particular, the psychology of deductive inference reveals human propositional reasoning to be vulnerable to logically-irrelevant linguistic effects, like co-references between premises [6]. Mental model theory provides an explanatory paradigm for these results [19], though prior work on the Erotetic Theory of Reasoning matches and has replicated mental model predictions on deductive inference tasks [20, 21].
In the language model literature, prior studies find that while modern LLMs largely succeed at syllogistic inference [22, 23, 24], they exhibit human-like reasoning failures, including distractability [25], content effects [26], and order effects [23, 27]. Closest to our work, the first application of the Erotetic Theory to LLMs found the predictive utility of the theory to increase for larger model sizes [28], though this study was limited to OpenAI’s GPT model family [29, 30, 31] and did not make use of PyETR. Our implementation yields an endlessly regenerative and content-agnostic test of the Erotetic Theory for LLMs in a manner more resilient to data contamination.
@@ 2 The Erotetic Theory of Reason
Modern accounts of human reasoning highlight a duality: impressive competence and predictable errors. The Erotetic Theory of Reasoning (ETR) explains both with one idea: we reason by managing questions and candidate answers (“alternatives”) and filtering them when new information arrives. Filtering is efficient but can drop relevant alternatives too early, producing characteristic fallacies. ETR formalises this process and defines when conclusions stabilise against follow-up questions (“erotetic equilibrium”).
At a high level (full formalism in Appendix A), ETR has three ingredients: (1) maintain disjunctive alternatives as candidate answers to an implicit question, (2) filter by best match with incoming information (risking premature elimination), and (3) recover by asking questions (raising structured alternatives) that reintroduce necessary information. This predicts a range of empirical reasoning data across propositional and first-order logic, probability, and decision-making. Morover, it is mathematically proved in [10] that ETR converges to normative standards of rationality in the limit, as the recovery step introduces enough alternatives to stabilise further reasoning steps.
Let us demonstrate ETR with a compact example (Example 49 in [10]).
@@
The correct logical answer is “not necessarily”, yet many respondents endorse “Mary has an ace” (60% in data following [11]).
To model the problem with logic, we would first choose a representation of the basic predicates, in this case $\mathsf{Ace}(x)$ for the predicate that (the card) $x$ is an ace, and $\mathsf{Has}(y,z)$ for the predicate that (player) $y$ holds (a card) $z$. The same applies when using ETR, where instead of logical formulas we have the central concept of view (see Appendix A) for the full mathematical formalism). It will suffice for now to use the (still precise) shorthand notation to express the premises and conclusion as ETR views: [^1] An additional aspect not present in ordinary logic is issue structure. This is notated with an asterisk on some occurrences of terms, indicative of being ‘at issue’ for the containing atomic proposition. The issue structure has no logical content, but plays a role in guiding ETR inferences. Determination of issue structure from cues in natural language or context is outside the scope of ETR, but a simple heuristic in this example is that the repetition of ‘Ace’ and ‘King’ suggests those concepts are at issue.
@@
Each view is given as a disjunctive set of alternatives, separated by a comma. Each alternative is to be considered as a conjunctive set of atomic propositions, where for convenience we omit the set-brackets and commas and simply juxtapose elements. Many readers will recognise these as disjunctive normal forms, but in ETR there is no inherent ordering to the atoms in a conjunction or alternatives in the disjunction. The individual lowercase letters can be understood are as existentially quantified variables, with the scope of the quantification encompassing the entire disjunction (‘prenex form’). A view may also carry a supposition, written as a superscript (see Table 1 for examples). Note that since any formula of first-order logic can be written in prenex disjunctive normal form, ETR is as expressive as full first-order logic.
PyETR, an open-source software package that serves as a calculator for ETR, can parse views from a similar text format.
@@
Note the requirement to explicitly declare existentially quantified names (the letter E may be used in place of $\exists$). ETR posits some default procedures for ordinary reasoning, built out of some basic operations. The workhorse among the basic operations is Update, which update a current view with an incoming one. We can check in PyETR that ETR does predict a fallacy here, because Update discards the seemingly less relevant alternative where John has a queen.
@@
>>> default_procedure_does_it_follow([p1,p2],$c$)
@@ 3 Methodology
In our experiment, a reasoning problem is given by a list of views (the premises), and the problem is to answer the question “what if anything follows?”. Our method generates reasoning problems where the ETR-predicted answer is a logical fallacy.
@@ 3.1 Data processing
We source our original bank of reasoning problems from examples presented in the Reason and Inquiry text [10]. These include basic “template” reasoning problems like modus ponens, modus tollens, and the disjunction fallacy as originally presented in Walsh and Johnson-Laird [6]. Table 1 presents the complete list of original problems. Not all reasoning problems take the form of valid syllogistic inferences; that is, not all conclusions from Erotetic Theory inference deductively follow from the premises.
@@ 3.2 Generative pipeline
Table 2: Logical Form Mutations
@@
Generating a new problem proceeds by iterative building a list of views (the premises in the problem). A new view is selected at random from the original problem bank and subjected to a random number of random mutation rules, it is added to the list if ETR predicts the problem admits a non-trivial answer. New views are added until all of the following stopping conditions are met. (1) The problem is the right size, meaning that summing the number of atoms in each view gives between 4 and 11. If 11 is exceeded, backtracking is used. (2) The ETR-predicted conclusion contains a single alternative (no disjunctions). (3) The ETR-predicted conclusion is a logical fallacy. To mitigate concerns about dataset-specific effects, we intentionally generated problems across diverse domains and structures. The PyETR-based generation ensures our problems test reasoning capacities rather than memorization of training examples.
@@
We initially generated 400 problems. Pre-analysis integrity checks identified 17 items whose fallacy status did not meet the final specification, so we excluded them prior to all analyses, yielding 383 problems. Given the modest reduction in sample size and the cost of re-running all model evaluations, we report results on this vetted 383-set; the generation pipeline supports regenerating larger sets for future work.
@@ 3.3 Mapping to natural language
To evaluate logical reasoning in language models in a more naturalistic context, we developed a systematic approach to convert formal logical statements into natural language prompts. We created 12 themes for natural language framings of logical problems, such as a researcher figuring out the properties of a novel element or a newly discovered creature. For each theme, we established consistent mappings between logical elements and thematic equivalents. The thematic conversion process serves multiple purposes: it ensures our results are robust against content effects [26]; it enables us to test whether the reasoning patterns are consistent in diverse contexts; and it mitigates potential data contamination, a significant problem for LLM evaluation [32, 33, 34, 35] by reformulating problems in novel scenarios unlikely to appear in training data.
Consider the following example of a disjunction fallacy problem in its formal representation: $\mathbf{P1}:\{Q(x)R(y),S(x)T(y)\},\mathbf{P2}:\{Q(x)\}$. In the “alchemy” theme, logical predicates (e.g., $A(x)$, $B(y)$) were mapped to thematic attributes (e.g., “is transmuting”, “is time-bending”), while logical variables (e.g., $x$, $y$) were mapped to imaginary thematic entities (e.g., “cosmic dust”, “vital mercury”). Using our alchemy theme, this was transformed into:
Im an alchemist studying mysterious substances in my laboratory. I need to understand their properties through logical reasoning. Heres what Ive discovered:
Cosmic dust is transmuting and vital mercury is time-bending, or cosmic dust is immortality-granting and vital mercury is spirit-affecting.
Also, we know that cosmic dust is immortality-granting.
What if anything follows? I do not have an intended answer in mind, and it is possible that nothing follows. Please be succinct and precise.
This is typical of the prompts sent to the LLM. It includes a thematic preamble and a natural language prompt to reason in an open-ended fashion with: “What if anything follows?” Each problem maintained the same prompt structure: an introduction establishing the thematic context, followed by the premises expressed in natural language, and concluding with a standardised question asking what follows from the premises. The prompt explicitly stated that nothing might follow, to avoid biasing models toward making unwarranted inferences [27].
@@ 3.4 Data Collection
Model Selection and Testing Conditions
We evaluated 38 language models of varying sophistication, including several specifically designed as reasoning models. Selection informally prioritized quantity, popularity of usage, and representation of a diverse range of scores. All models were accessed via OpenRouter and tested using LLMharness. Chatbot Arena scores were obtained from https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard/resolve/main/elo_results_20250423.pkl. Scores for the HELM Capabilities benchmark for 9 models were obtained from https://crfm.stanford.edu/helm/capabilities/v1.10.0/#/leaderboard [17] and estimates for training compute for 19 models were obtained from https://epoch.ai/data/ai-models [16] (both URLs accessed 30 July 2025).
To ensure a fair comparison across architectural differences, all reasoning models were allocated 2,400 thinking tokens during evaluation. This standardization was justified given our use of the Chatbot Arena Elo score as our primary capability metric, which already accounts for performance differences regardless of underlying mechanisms [14].
@@
Models provided conclusions in natural language, which we converted to PyETR format for evaluation. To ensure fair assessment of reasoning (not formatting), GPT-4.1-mini served as our translation layer without access to the original premises. Spot-checking confirmed translation fidelity.
@@
The full set of logical problems were framed as evals, using Eleuther’s Language Model Evaluation Harness [36], which handles evaling and metric collection against a wide range of models. OpenRouter [37] was used for all models, in order to have a unified API. This consumed less than \$1,000 of compute resources. Each model was given a token limit of 3,000 output tokens.
@@
We collected the following measures to evaluate both the logical correctness of model responses and their alignment with human-like reasoning patterns, enabling our analysis of how these patterns correlate with model sophistication. A model’s answer was considered ‘logically correct’ if it was a logical consequence of the premises. This was tested using PySMT [38] to check whether the negation of the conclusion was inconsistent with the premises. LLM responses were classified as ‘ETR-predicted’ if ETR predicts an endorsement of that conclusion according to PyETR’s implementation of default_procedure_does_it_follow.
@@
We selected Pearson correlation to capture linear relationships between variables while also reporting Spearman’s rank correlation to account for potentially non-linear monotonic relationships without assuming normality in the distribution of fallacy rates. For our premise-order intervention analysis, we employed two-proportion z-tests to rigorously evaluate whether reversing premise order produced statistically significant changes in fallacy rates. The z-test was specifically chosen because it allows us to determine if the observed differences in proportions (original vs. reversed premise order) represent genuine effects rather than random variation, providing a reliable measure of the significance of order effects across different model capabilities. This statistical approach enables us to quantify the extent to which models exhibit the same premise-order sensitivity documented in human reasoning studies.

### v3

@@ 1 Introduction
Language models increasingly solve complex tasks (Wei et al., 2023; Kojima et al., 2022). We ask a simple question: when they err on controlled reasoning problems, do their errors align with human fallacies? Human reasoning shows systematic, repeatable fallacies (Tversky and Kahneman, 1974; Kahneman and Tversky, 1982; Evans, 1994; Walsh and Johnson-Laird, 2004; Johnson-Laird, 2006). These are not random mistakes; they follow predictable patterns across tasks. Understanding whether LLM errors share these patterns is useful both scientifically and for deployment in settings that require reliable reasoning (Jacobs, 2021; Bommasani et al., 2022).
To rigorously investigate this question, we leverage the Erotetic Theory of Reasoning (ETR) (Koralus, 2022; Mascarenhas and Koralus, 2017a; Koralus and Mascarenhas, 2013), a formal cognitive theory that precisely predicts human reasoning patterns across multiple domains. ETR posits that human reasoning operates by maintaining disjunctive alternatives and filtering these alternatives when taking on new information, a process that can systematically lead to error. ETR provides formal specifications of both how and when humans will make specific reasoning errors, allowing us to generate arbitrary reasoning problems with predictable failure patterns.
ETR has an open-source implementation, PyETR (Koralus et al., 2025). Using PyETR, we mechanically generate reasoning problems predicted to elicit specific fallacies. We evaluate 38 models on a fixed corpus of 383 such problems.
Our experimental investigation evaluates 38 language models ranging from smaller models (e.g., Mistral 7B Instruct v0.1) to larger systems (e.g., GPT-4.5, Claude 3.7), treating Chatbot Arena Elo as one capability proxy (Chiang et al., 2024). We observe that as this proxy increases, a larger share of logically incorrect answers align with ETR-predicted human fallacies ($\rho=0.360,p=0.0265$). This is a correlative claim; our experiments do not suggest or support any causal mechanism.
While language models often improve on many benchmarks (Wei et al., 2023; Bubeck et al., 2023), our results indicate that within our task domain, the proportion of errors matching ETR-predicted fallacies increases with Chatbot Arena Elo as a capability proxy. These findings suggest an overlap between LLM error patterns and human fallacy patterns irrespective of any underlying cognitive mechanisms.
We affirm the robustness of our result using metrics other than Chatbot Arena Elo, where available, for a subset of the 38 models. We find significant Spearman correlation with estimates of training compute (Epoch AI, 2025) ($\rho=0.489,p=0.0334$) and significant exponential fit for mean score on the HELM capabilities benchmark (Liang et al., 2023) ($r=0.796,p=0.0103$).
@@
The remainder of the paper is organised as follows. Section 2 introduces ETR and PyETR. Section 3 details methodology. Section 4 presents results. Section 5 discusses implications and concludes.
@@
Substantial prior literature in psychology and cognitive science documents irrationalities in human judgement and decision-making (Tversky and Kahneman, 1974; Kahneman and Tversky, 1982; Evans, 1994; Kahneman et al., 1990). In particular, the psychology of deductive inference reveals human propositional reasoning to be vulnerable to logically-irrelevant linguistic effects, like co-references between premises (Walsh and Johnson-Laird, 2004). Mental model theory provides an explanatory paradigm for these results (Johnson-Laird and Byrne, 1991), though prior work on the Erotetic Theory of Reasoning matches and has replicated mental model predictions on deductive inference tasks (Mascarenhas and Koralus, 2015; 2017b).
In the language model literature, prior studies find that while modern LLMs largely succeed at syllogistic inference (Clark et al., 2021; Eisape et al., 2024; Bertolazzi et al., 2024), they exhibit human-like reasoning failures, including distractability (Shi et al., 2023), content effects (Lampinen et al., 2024), and order effects (Eisape et al., 2024; Saparov and He, 2023). In a prior work by two of the present authors, the first application of the Erotetic Theory to LLMs found the predictive utility of the theory to increase for larger model sizes (Koralus and Wang-Maścianica, 2023), though this study was limited to OpenAI’s GPT model family (Brown et al., 2020; OpenAI, 2022; 2024) and did not make use of PyETR. Our implementation yields an endlessly regenerative and content-agnostic test of the Erotetic Theory for LLMs in a manner more resilient to data contamination.
@@ 2 The Erotetic Theory of Reason
Modern accounts of human reasoning highlight a duality: impressive competence and predictable errors. The Erotetic Theory of Reasoning (ETR) explains both with a single idea: we reason by managing questions and candidate answers (or “alternatives”) and filtering them when new information arrives. Filtering is efficient but can drop relevant alternatives too early, producing characteristic fallacies. ETR formalises this process and defines when conclusions stabilise against follow-up questions (a state called “erotetic equilibrium”).
At a high level, [^1] A full formalism for ETR is provided in Appendix A. ETR considers reasoning to consist of three steps: (1) maintain disjunctive alternatives as candidate answers to an implicit question, (2) filter by best match with incoming information (risking premature elimination), and (3) recover by asking questions (raising structured alternatives) that reintroduce necessary information. This theory predicts a range of empirical reasoning data across propositional and first-order logic, probability, and decision-making. Morover, it is mathematically proved in Koralus (2022) that ETR converges to normative standards of rationality in the limit, as the recovery step introduces enough alternatives to stabilise further reasoning steps.
We can demonstrate ETR with a compact example (Example 49 in Koralus (2022)).
@@
The correct logical answer is “not necessarily”, yet many respondents endorse “Mary has an ace” (60% in data following Mascarenhas and Koralus (2017a)).
To model the problem in logic, we would first choose a representation of the basic predicates, in this case $\mathsf{Ace}(x)$ for the predicate that (the card) $x$ is an ace, and $\mathsf{Has}(y,z)$ for the predicate that (player) $y$ holds (a card) $z$. The same applies in ETR, where instead of logical formulas we have the central concept of a view (see Appendix A) for the full mathematical formalism). It will suffice for now to use the (still precise) shorthand notation to express the premises and conclusion as ETR views: [^2] An additional aspect not present in ordinary logic is issue structure. This is notated with an asterisk on some occurrences of terms, indicative of being ‘at issue’ for the containing atomic proposition. The issue structure has no logical content, but plays a role in guiding ETR inferences. Determination of issue structure from cues in natural language or context is outside the scope of ETR, but a simple heuristic in this example is that the repetition of ‘Ace’ and ‘King’ suggests those concepts are at issue.
@@
Each view is given as a disjunctive set of alternatives, separated by a comma. Each alternative is to be considered as a conjunctive set of atomic propositions, where for convenience we omit the set-brackets and commas and simply juxtapose elements. Many readers will recognise these as disjunctive normal forms, but in ETR there is no inherent ordering to the atoms in a conjunction or alternatives in the disjunction. The individual lowercase letters can be understood as existentially quantified variables, with the scope of the quantification encompassing the entire disjunction (“prenex form”). A view may also carry a supposition, written as a superscript (see Table 1 for examples). Note that since any formula of first-order logic can be written in prenex disjunctive normal form, ETR is as expressive as full first-order logic.
PyETR, an open-source software package that serves as a calculator for ETR (Koralus et al., 2025), can parse views from a similar text format.
@@
Note the requirement to explicitly declare existentially quantified names (the letter E may be used in place of $\exists$). ETR posits some default procedures for ordinary reasoning built out of basic operations. The workhorse among the basic operations is Update, which updates a current view with an incoming one. We can check in PyETR that ETR does predict a fallacy here, because Update discards the seemingly less relevant alternative where John has a queen.
@@
>>> default_procedure_does_it_follow([p1,p2],q)
@@ 3 Methodology
In our experiment, a reasoning problem is given by a list of views (the premises), and the problem is to answer the question “what, if anything, follows?”. Our method generates reasoning problems where the ETR-predicted answer is a logical fallacy.
@@ 3.1 Data processing
We source our original bank of reasoning problems from examples presented in the Reason and Inquiry text (Koralus, 2022). These include basic “template” reasoning problems like modus ponens, modus tollens, and the disjunction fallacy as originally presented in Walsh and Johnson-Laird (2004). Table 1 presents the complete list of original problems. Not all reasoning problems take the form of valid syllogistic inferences; that is, not all conclusions from ETR inference deductively follow from the premises.
@@ 3.2 Generative pipeline
Table 2: Mutation rules for ETR views.
@@
Generating a new problem proceeds by iteratively building a list of views (the premises in the problem). A new view is selected at random from the original problem bank and subjected to a random number of random mutation rules. The mutated view is added to the list if ETR predicts the problem admits a non-trivial answer. New views are added until all of the following stopping conditions are met: (1) The problem is the right size, meaning that summing the number of atoms in each view gives between 4 and 11. If 11 is exceeded, backtracking is used. (2) The ETR-predicted conclusion contains a single categorical alternative (without disjunctions). (3) The ETR-predicted conclusion is a logical fallacy. To mitigate concerns about dataset-specific effects, we intentionally generated problems across diverse domains and structures. The PyETR-based generation ensures our problems test reasoning capacities rather than memorization of training examples.
@@
We initially generated 400 problems. Pre-analysis integrity checks identified 17 items whose fallacy status did not meet the above stopping conditions, so we excluded them prior to all analyses, yielding 383 problems. Given the modest reduction in sample size and the cost of re-running all model evaluations, we report results on this vetted 383-set; the generation pipeline supports regenerating larger sets for future work.
@@ 3.3 Mapping to natural language
To evaluate logical reasoning in language models in a more naturalistic context, we developed a systematic approach to convert formal logical statements into natural language prompts. We created 12 themes for natural language framings of logical problems, such as a researcher figuring out the properties of a novel element or a newly discovered creature. The elements of our 12 themes are listed in Appendix B. For each theme, we established consistent mappings between logical elements and thematic equivalents. The thematic conversion process serves multiple purposes: it ensures our results are robust against content effects (Lampinen et al., 2024), it enables us to test whether the reasoning patterns are consistent in diverse contexts, and it mitigates potential data contamination, a significant problem for LLM evaluation (Zhou et al., 2023; Deng et al., 2024; Li and Flanigan, 2024; Singh et al., 2024) by reformulating problems in novel scenarios unlikely to appear in training data.
Consider the following example of a disjunction fallacy problem in its formal representation:
P1 $\{Q(x)R(y),S(x)T(y)\}$
P2 $\{Q(x)\}$
In the “alchemy” theme, logical predicates (e.g., $Q(x)$, $R(y)$) were mapped to thematic attributes (e.g., “is transmuting”, “is time-bending”), while logical variables (e.g., $x$, $y$) were mapped to imaginary thematic entities (e.g., “cosmic dust”, “vital mercury”). Using our alchemy theme, this was transformed into:
I’m an alchemist studying mysterious substances in my laboratory. I need to understand their properties through logical reasoning. Here’s what I’ve discovered: Cosmic dust is transmuting and vital mercury is time-bending, or cosmic dust is immortality-granting and vital mercury is spirit-affecting. Also, we know that cosmic dust is immortality-granting. What if anything follows? I do not have an intended answer in mind, and it is possible that nothing follows. Please be succinct and precise.
This is typical of the prompts sent to the LLM. [^3] Complete prompts, with reasoning traces and responses, are provided in Appendix D. It includes a thematic preamble and a natural language prompt to reason in an open-ended fashion with: “What if anything follows?” Each problem maintained the same prompt structure: an introduction establishing the thematic context, followed by the premises expressed in natural language, and concluding with a standardised question asking what follows from the premises. The prompt explicitly stated that nothing might follow, to avoid biasing models toward making unwarranted inferences (Saparov and He, 2023).
@@ 3.4 Data Collection
Model Selection
We evaluated 38 language models of varying sophistication, including several specifically designed as reasoning models. Selection informally prioritized quantity, popularity of usage, and representation of a diverse range of scores. Chatbot Arena scores were obtained from https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard/resolve/main/elo_results_20250423.pkl. Scores for the HELM Capabilities benchmark for 9 models were obtained from https://crfm.stanford.edu/helm/capabilities/v1.10.0/#/leaderboard (Liang et al., 2023) and estimates for training compute for 19 models were obtained from https://epoch.ai/data/ai-models (Epoch AI, 2025) (both URLs accessed 30 July 2025).
@@
Models provided conclusions in natural language, which we converted to PyETR format for evaluation. To ensure fair assessment of reasoning (not formatting), GPT-4.1-mini served as our translation layer without access to the original premises. Our manual spot-checks confirmed this translation step to be perfectly faithful.
@@
The full set of logical problems were framed as evals, using Eleuther’s Language Model Evaluation Harness (Gao et al., 2024), which handles evaling and metric collection against a wide range of models. OpenRouter (OpenRouter, 2025) was used for all models in order to have a unified API. This consumed less than \$1,000 of compute resources. Each model was given a token limit of 3,000 output tokens.
To ensure a fair comparison across architectural differences, all reasoning models were allocated 2,400 thinking tokens during evaluation. This standardization was justified given our use of the Chatbot Arena Elo score as our primary capability metric, which already accounts for performance differences regardless of underlying mechanisms (Chiang et al., 2024).
@@
We collected the following measures to evaluate both the logical correctness of model responses and their alignment with human-like reasoning patterns, enabling our analysis of how these patterns correlate with model sophistication. A model’s answer was considered ‘logically correct’ if it was a logical consequence of the premises. This was tested using PySMT (Cimatti et al., 2017) to check whether the negation of the conclusion was inconsistent with the premises. LLM responses were classified as ‘ETR-predicted’ if ETR predicts an endorsement of that conclusion according to PyETR’s implementation of default_procedure_does_it_follow.
@@
We selected Pearson correlation to capture linear relationships between variables while also reporting Spearman’s rank correlation to account for potentially non-linear monotonic relationships without assuming normality in the distribution of fallacy rates. For our premise-order intervention analysis (discussed in Section 4.4), we employed two-proportion z-tests to rigorously evaluate whether reversing premise order produced statistically significant changes in fallacy rates. The z-test was specifically chosen because it allows us to determine if the observed differences in proportions (original vs. reversed premise order) represent genuine effects rather than random variation, providing a reliable measure of the significance of order effects across different model capabilities. This statistical approach enables us to quantify the extent to which models exhibit the same premise-order sensitivity documented in human reasoning studies.